# Quantitative and Predictive Folding Models from Limited Single-Molecule Data Using Simulation-Based Inference

### v1

@@ Abstract
The study of biomolecular folding has been greatly advanced by single-molecule force spectroscopy (SMFS), which enables the observation of the dynamics of individual molecules. However, extracting quantitative models of fundamental properties such as folding landscapes from SNFS data is very challenging due to instrumental noise, linker artifacts, and the inherent stochasticity of the process, often requiring extensive datasets and complex calibration experiments. Here, we introduce a framework based on simulation-based inference (SBI) that overcomes these limitations by integrating physics-based modeling with deep learning. We apply this framework to analyze constant-force measurements of a DNA hairpin. From a single, short experimental trajectory of only two seconds, we successfully reconstruct the hairpin’s free energy landscape and folding dynamics, obtaining results that are in close agreement with established deconvolution methods that require approximately 100 times more data. Furthermore, the Bayesian nature of this approach robustly quantifies uncertainties for inferred parameter values, including the free-energy profile, diffusion coefficients, and linker stiffness, without needing independent measurements of instrumental properties. The inferred model is predictive, generating simulated trajectories that quantitatively reproduce the thermodynamic and kinetic properties of the experimental data. This work establishes SBI as a highly efficient and powerful tool for analyzing single-molecule experiments. The ability to derive statistically robust models from minimal datasets is crucial for investigating complex biomolecular systems where extensive data collection is impractical or impossible. Consequently, our SBI framework enables the rigorous quantitative analysis of previously intractable biomolecular systems, paving the way for novel applications of SMFS.
@@
Introduction - Biomolecular folding is a fundamental process in which proteins, nucleic acids, and other biopolymers adopt three-dimensional structures essential for their function [1]. Understanding these complex folding mechanisms is crucial to elucidating the fundamental principles of biomolecular self-organization and its connection to diseases. The folding dynamics is governed by a complex interplay between many intermolecular forces, often modeled as a diffusive process on a rugged free energy landscape [2]. Characterization of the folding free energy landscape provides a complete thermodynamic picture of the folding process, providing crucial insights into folding mechanisms and kinetics. However, experimentally obtaining free energy landscapes is exceptionally challenging due to fundamental limitations in spatial and temporal resolution [3].
Single-molecule force spectroscopy (SMFS) probes the folding and unfolding dynamics of individual biomolecules held under tension [4, 5]. These measurements allow the reconstruction of free energy landscapes as a function of a one-dimensional order parameter, usually the molecular extension [6, 7, 8, 9, 10]. Free energy landscapes have been reconstructed from constant-force [11], constant-trap [12, 13], and non-equilibrium measurements [14].
However, SMFS experiments use a large pulling device, which is attached to the biomolecule via long and flexible linkers. The measured extension is therefore only an indirect proxy of the molecular one, and is a convolution of the intrinsic molecular dynamics with the dynamics of the linker and the slow response of the pulling device [15, 16]. The measurement noise and the inherent stochasticity of single-molecule trajectories further complicate the estimation of the underlying free energy landscape. In principle, it is possible to deconvolve the linker contribution from the measured trajectory to estimate the hidden molecular extension and, therefore, the free energy landscape [17, 7, 11]. In practice, this requires large amounts of data and a very precise characterization of the linker, which is technically challenging and very labor-intensive. Consequently, the full profile of the folding landscape has been reconstructed for only a handful of molecules.
@@
An alternative approach that can overcome these problems is to reconstruct the free energy landscape by fitting a dynamical model $\mathcal{M}(\bm{\theta})$ to the observed trajectories $\bm{q}_{[1:N]}$, a vector containing $N$ measured extensions sampled at equal time intervals. The parameters $\bm{\theta}$ determine the landscape, and must be inferred from the observed trajectory [18]. Bayesian inference provides a systematic statistical framework to determine the posterior distribution of model parameters $\bm{\theta}$ that is most consistent with $\bm{q}_{[1:N]}$:
$p(\bm{\theta}|\bm{q}_{[1:N]})=\frac{p(\bm{q}_{[1:N]}|\bm{\theta})p(\bm{\theta})}{p(\bm{q}_{[1:N]})}$ (1)
@@
Here, $p(\bm{\theta})$ is the prior, encoding all knowledge before the inference. The likelihood $p(\bm{q}_{[1:N]}|\bm{\theta})$, is the probability of observing the trajectory $\bm{q}_{[1:N]}$ given parameters $\bm{\theta}$. $p(\bm{q}_{[1:N]})$ is a normalization factor.
@@
Typical inference algorithms require maximizing the likelihood with respect to the parameters. However, evaluating the likelihood for fitting only partially observable dynamical models is very challenging, because it is a marginalized likelihood over all possible “latent” trajectories of the actual molecular extension $\bm{x}_{[1:N]}$ [19, 20]:
$p(\bm{q}_{[1:N]}|\bm{\theta})=\int\mathcal{D}\bm{x}_{[1:N]}p(\bm{q}_{[1:N]},\bm{x}_{[1:N]}|\bm{\theta})$ (2)
@@
An emerging approach to tackle this problem is simulation-based inference (SBI) [21, 18, 20], which integrates physics-based models with probabilistic deep learning to learn a surrogate model of the posterior distribution [22], likelihood [23], or likelihood ratio [24]. The core idea is to use a simulator to generate synthetic measurements for specific parameters. Using deep neural networks, we can leverage the simulations to learn a surrogate model of the posterior distribution. SBI has shown success across various fields, from astrophysics [25, 26] to neuroscience [27, 28], effectively tackling challenges associated with complex likelihood functions [29].
@@
Here, we showcase how to leverage SBI to learn quantitative folding models from SMFS experiments. Using a well-characterized 30R50/T4 DNA hairpin as a model system [30, 31, 32, 16], we reconstruct the free energy landscape from limited experimental data in a straightforward way. Taking advantage of the Bayesian framework, we also obtain accurate uncertainties of our estimates and quantitative predictive models. This work establishes an efficient and robust method for quantitative analysis of SMFS experiments.
@@
Results - Single nucleic acid hairpins serve as an ideal model system for studying biomolecular folding dynamics [33]. These self-complementary sequences fold into simple structures, offering a well-defined experimental framework for investigating folding mechanisms. The end-to-end distance of the DNA serves as a convenient reaction coordinate for the folding reaction, along which the dynamics are well characterized by one-dimensional Brownian diffusion on an effective free energy landscape [30, 34]. Beyond their use as model systems for biomolecular folding, DNA hairpins also have biological relevance as they play important roles in a multitude of biological processes [35, 36].
@@
In this study, we apply our inference framework [18] for the first time to learn a quantitative folding model from constant-force measurements of a 30R50/T4 DNA hairpin. This particular hairpin’s folding mechanism has been extensively studied, with measurements of the folding dynamics spanning from microseconds to minutes and folding energy landscape reconstructions using various methods and experimental setups [7, 37, 34].
@@
We start with a physical model $\mathcal{M}(\bm{\theta})$, parametrized by $\bm{\theta}$, which describes both the biomolecular folding process and the measurement of the observed data. Given a measured trajectory $\bm{q}_{[1:N]}$, we aim to determine the optimal parameter set $\bm{\theta}$ that best fits the measurement by computing the posterior $p(\bm{\theta}|\bm{q}_{[1:N]}$). According to Bayes’ theorem, the posterior for this model is given by Eq. 1.
Figure 2: Free energy profile reconstruction. (A) Experimental time series used for inference. (B) Reconstructed free energy profile. Best estimate $\hat{\bm{\theta}}^{\textrm{exp}}_{\textrm{MAP}}$ (Maximum a posteriori, MAP) in red, and posterior samples covering a 68 % confidence interval as blue thin lines. The black line indicates the estimate using deconvolution. (C) Best free energy profile (MAP) estimate for 20 independent experimental time series.
@@
The harmonic-spring model $\mathcal{M}(\bm{\theta})$ provides a well-established framework for describing single-molecule force spectroscopy experiments [19, 15, 10]. This model characterizes the dynamics of the combined system of biomolecule and measuring apparatus as a diffusive process on a two-dimensional free energy surface $G(q,x)$. Here, $q$ is the measured extension, which includes the extension of both the linker and the molecule, whereas $x$ is the (hidden) molecular extension. The surface can be decomposed into two parts: $G(q,x)=G_{0}(x)+1/2\cdot k_{l}(q-x)^{2}$ where $G_{0}(x)$ is the intrinsic free energy landscape of the molecular folding reaction that we wish to characterize, while the second term describes the linker coupling the molecule with the apparatus. We assume anisotropic Brownian dynamics on this effective energy landscape, with diffusion coefficient $D_{x}$ for the molecule and $D_{q}$ for the apparatus. The molecular free energy profile $G_{0}(x)$ is modeled using cubic spline interpolation. The model parameters $\bm{\theta}=\{D_{q}/D_{x},k_{l},G_{0}(x_{0}),G_{0}(x_{1}),...,G_{0}(x_{N_{\textrm{node}}})\}$ are the ratio of diffusion coefficients $D_{q}/D_{x}$, the linker stiffness $k_{l}$, and $N_{\textrm{node}}$ positions of the spline nodes $G_{0}(x_{i})$ to characterize the free energy profile $G_{0}(x)$. By inferring $\bm{\theta}$ from a measured trajectory $\bm{q}_{[1:N]}$, we can fully characterize the biomolecular folding as a diffusive process on the 1-dimensional energy landscape $G_{0}(x)$.
@@
Direct computation of the likelihood (Eq. 2) is extremely challenging because it requires integrating over all hidden molecular trajectories $\bm{x}_{[1:N]}$. In other words, to compute the likelihood, we have to search over all possible paths $\bm{x}_{[1:N]}$ the molecule could have taken while observing $\bm{q}_{[1:N]}$. In contrast, simulating a measured trajectory $\bm{q}_{[1:N]}\sim p(\bm{q}_{[1:N]}|\bm{\theta})$ is straightforward. This is the key idea behind SBI, a novel approach for Bayesian inference, which combines physics-based modeling and deep learning in order to approximate the posterior $p(\bm{\theta}|\bm{q}_{[1:N]})$ (Fig. 1) from simulated data. SBI offers a general solution to perform Bayesian inference in settings where the likelihood is either intractable or too expensive to evaluate explicitly. SBI starts by generating prior samples $\bm{\theta}_{i}\sim p(\bm{\theta})$, which are used to simulate synthetic trajectories $\bm{q}_{[1:N],i}\sim\mathcal{M}(\bm{\theta}_{i})$ (Fig. 1 upper panel). The simulated data set $\mathcal{D}=\{(\bm{q}_{[1:N],i},\bm{\theta}_{i})\}_{i=1}^{N}$ is then subsequently used to train a flexible neural network-based density estimator $f_{\psi}(\bm{\theta}|\bm{q}_{[1:N]})$ of parameters $\psi$ to approximate the Bayesian posterior (Fig. 1 center panel). To learn the posterior from $\mathcal{D}$, the neural network parameters $\psi$ are optimized by minimizing the loss function $\mathcal{L}(\psi)=1/N\sum_{i}^{N}f_{\psi}(\bm{\theta}_{i}|\bm{q}_{[1:N],i})$ on $\mathcal{D}$. After training, the posterior can be evaluated with an experimental trajectory via $p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})\approx f_{\psi}(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$, allowing us to infer the underlying folding parameters (Fig. 1 right panel).
Figure 3: Diffusion coefficients and linker stiffness estimates. Posteriors obtained using 20 independent experimental time series, quantifying the inference on (A) the ratio of diffusion coefficients $D_{q}/D_{x}$, and (B) the linker stiffness $k_{l}$.
@@
We analyzed data from the 30R50/T4 DNA hairpin (Fig. 2A Cartoon). Our analysis used a short 2-second long constant force measurement $\bm{q}_{[1:N]}^{\textrm{exp}}$ (Fig. 2 A) that captured approximately seven transitions between the folded and unfolded states. Our objective was to extract a folding model that accurately describes the observed folding dynamics in our experimental data $\bm{q}_{[1:N]}^{\textrm{exp}}$.
@@
We performed Sequential Neural Posterior Estimation (SNPE) to infer the posterior distribution of model parameters $p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$ from the experimental trajectory $\bm{q}_{[1:N]}^{\textrm{exp}}$. SNPE iteratively approximates the posterior distribution through alternating rounds of simulation and posterior density estimation. In each round, we performed a batch of simulations using parameters drawn from the current posterior estimate, which then served as the prior for the following round. After 14 rounds, each involving 2000 simulations, we obtained a surrogate posterior $p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$ (SI Figure 1) in about 11 hours (on one Xeon Skylake Gold 6148 CPU). After training, we can evaluate the surrogate posterior on experimental trajectories. We can either sample parameters according to the posterior distribution $\bm{\theta}^{\textrm{exp}}\sim p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$, exploring many parameters which explain the experimental data. Alternatively, we can also identify the maximum a posteriori (MAP) estimate $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}=\textrm{argmax}_{\bm{\theta}}\ p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$, the parameter which best explains the data.
@@
To validate our results, we compared our inferred free energy landscape with one obtained through deconvolution (Figure 2B black dotted line). Deconvolution is an established method that reconstructs the free energy landscape by iteratively removing measurement noise from the experimental data (see SI)[7]. To obtain the biomolecule free energy profile $G_{0}(x)$ with deconvolution, two measurements are performed: one of the biomolecule attached to the linker yielding $G(q)$, and another of the linker alone. By measuring the linker contribution $V(q-x)$ separately, its artifacts can be removed from the observed data. Deconvolution aims to find the free energy profile $G_{0}(x)$ that, when convolved with the linker noise characteristics $V(q-x)$, reproduces the experimentally observed profile $G(q)$. While both approaches yield closely matching profiles, confirming the reliability of our inference, deconvolution requires substantially more experimental data—approximately 20-100 times as many folding/unfolding transitions, which makes the analysis much more labor intensive and prone to errors arising from instrumental drift. Furthermore, deconvolution requires a separate, very careful characterization of the linker, which, if done incorrectly, can lead to errors in the resulting free energy landscape.
Additionally, we demonstrated that our estimated uncertainties are accurate. We applied our inference scheme to 20 independent 2-second measurements $\{\bm{q}_{[1:N],i}^{\mathrm{exp}}\}_{i=1}^{N=20}$ of the same DNA hairpin molecule. The results, shown in Figure 2C, yield statistically consistent free energy landscapes across all measurements. Importantly, the deviations among these independently inferred profiles (red lines in Figure 2C) fall within the uncertainty bounds (blue lines in Figure 2B) determined from analyzing a single trajectory. This overlap indicates that the error model of the surrogate posterior is well-calibrated.
The model we fitted to the experimental data $\bm{q}_{[1:N]}^{\textrm{exp}}$ explains more than just the molecular free energy profile. We also obtained estimates for the ratio of diffusion coefficients $D_{q}/D_{x}$ and linker stiffness $k_{l}$. Figure 3 reports the marginal posterior distribution for $D_{q}/D_{x}$ and $k_{l}$ for 20 independent experimental trajectories. We also observed a systematic overlap of the marginal posterior distributions for these parameters. This overlap indicates again that our error model is accurate.
To further assess our model’s accuracy and predictive power, we simulated trajectories using the best-fitting parameters, i.e., $\bm{q}_{[1:N],i}\sim p(\bm{q}_{[1:N]}\mid\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},i})$. We then compared these simulated trajectories to the experimental trajectory used for parameter inference (Figure 4A). If the model is adequate and the fit accurate, we expect that the experimental $\bm{q}_{[1:N]}^{\mathrm{exp}}$ and simulated trajectory $\bm{q}_{[1:N]}$ look very similar.
@@
Figure 4: Predictive check using simulations with best-fitting parameters $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}$. (A) Trajectory $\bm{q}_{[1:N],i=1}$ simulated with $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},i=1}$ (blue) compared to experimental trajectory $\bm{q}_{[1:N],i=1}^{\mathrm{exp}}$ used to make the inference (red). (B) The potential of mean force estimated from 20 synthetic trajectories $\bm{q}_{[1:N],i}$ (blue) simulated with $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},i}$ compared to the potential of mean force obtained from the experimental trajectories (red). (C) Autocorrelation functions for segments of the trajectories in the folded and unfolded states, respectively, comparing experimental $\bm{q}_{[1:N],i}^{\mathrm{exp}}$ (red) and simulated trajectories $\bm{q}_{[1:N],i}$ (blue).
@@
While this visual comparison is only qualitative, it shows excellent agreement between the simulated and experimental trajectories. To make this comparison more quantitative, we created a simulated data set to mimic exactly our experimental one. We therefore performed a total of 20 simulations $\{\bm{q}_{[1:N],i}\}_{i=1}^{N=20}$ using the best fitting parameter $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},i}$ from each experimental trajectory. First, we compared the thermodynamics of the observed trajectories by inspecting the potential of mean force (PMF) $G(q)$. By combining all trajectories, we computed the average PMF $G(q)$ by binning them along $q$. Figure 4B compares the average PMF calculated with the simulated and the one with the experimental trajectories. While the PMFs from the simulations and the experiment align very well in both the unfolded and folded states, the PMFs from the simulated trajectories show a slightly higher energy barrier in the transition region. To asses the variability around the average PMF we computed the PMF for all the individual trajectories (SI Figure 3).
@@
We additionally compared the observed kinetics of the simulated and the experimental data by analyzing the transitions (folding and unfolding) in the observed trajectories (see SI). The experimental trajectories show transition rate of $2.8\pm 0.3\leavevmode\nobreak\ \mathrm{s^{-1}}$, and for the simulated trajectories, we estimated a transition rate of $2.2\pm 0.2\leavevmode\nobreak\ \mathrm{s^{-1}}$. This close agreement between the experimental and simulated transition rates demonstrates that our model successfully captures the underlying kinetics.

### v2

# arXiv:2508.02509v2[physics.chem-ph] 02 Mar 2026

@@ Abstract
The study of biomolecular folding has been greatly advanced by single-molecule force spectroscopy (SMFS), which enables the observation of the dynamics of individual molecules. However, extracting quantitative models of fundamental properties such as folding landscapes from SMFS data is very challenging due to instrumental noise, linker artifacts, and the inherent stochasticity of the process, often requiring extensive datasets and complex calibration. Here, we introduce a framework based on simulation-based inference (SBI) that overcomes these limitations by integrating physics-based modeling with deep learning. We first apply this framework to analyze constant-force measurements of a DNA hairpin. From a single experimental trajectory of only two seconds, we successfully reconstruct the hairpin’s free energy landscape and folding dynamics, obtaining results in close agreement with established deconvolution methods that require 10–100 times more data. Furthermore, we demonstrate the generality of our approach by applying it to a riboswitch aptamer featuring multiple states and tertiary contacts, resolving the profile of a landscape featuring four metastable states from a single trajectory. The Bayesian nature of this approach robustly quantifies uncertainties for all inferred parameters, including diffusion coefficients and linker stiffness, without needing independent measurements of instrument properties. The inferred models are predictive, generating simulated trajectories that quantitatively reproduce experimental thermodynamics and kinetics. The ability to derive statistically robust models from minimal datasets is crucial for investigating complex biomolecular systems where extensive data collection is impractical, paving the way for novel applications of SMFS.
@@
Biomolecular folding is a fundamental process in which proteins, nucleic acids, and other biopolymers adopt three-dimensional structures essential for their function [9]. Understanding folding mechanisms is crucial to elucidating the principles of biomolecular self-organization and its connection to diseases. The folding dynamics arise from a complex interplay between many intermolecular forces, often modeled as a diffusive process on a rugged free energy landscape [8]. Characterizing the folding free energy landscape provides a complete thermodynamic picture of the folding process, providing insights into mechanisms and kinetics. However, experimentally obtaining free energy landscapes remains challenging due to limitations in spatial and temporal resolution [2].
Single-molecule force spectroscopy (SMFS) probes the folding dynamics of individual biomolecules held under tension [42, 32]. These measurements enable the reconstruction of free energy landscapes along a one-dimensional reaction coordinate, usually the molecular extension [25, 53, 12, 13, 26]. Such reconstructions have been achieved using constant-force [52], constant-position [19, 51], and non-equilibrium measurements [22].
@@
However, SMFS measurements are indirect: the biomolecule is connected to a large pulling device through long, flexible linkers, so the measured extension reflects a convolution of molecular dynamics, linker fluctuations, and the instrument response [3, 36]. Measurement noise and the inherent stochasticity of single-molecule trajectories further complicate the estimation of the underlying free energy landscape. In principle, the linker contribution can be removed by deconvolution [23, 53, 52]. In practice, this requires a large amount of data and a very precise characterization of the linker, which is technically challenging and labor-intensive. Consequently, accurate free energy reconstructions exist only for a handful of molecules.
@@
An alternative approach to overcome these problems is to reconstruct the free energy landscape by fitting a dynamical model $\mathcal{M}(\bm{\theta})$ to the observed trajectory $\bm{q}_{[1:N]}$, a vector containing $N$ measured extensions sampled at equal time intervals. The parameters $\bm{\theta}$ determine the landscape, and must be inferred from the observed trajectory [10]. Bayesian inference provides a systematic statistical framework to determine the posterior distribution over $\bm{\theta}$ that is most consistent with $\bm{q}_{[1:N]}$:
{align}
@@
p(θ|q_[1:N]) = p(q[1:N]|θ)p(θ)p(q[1:N]) Here, $p(\bm{\theta})$ is the prior, encoding prior knowledge, $p(\bm{q}_{[1:N]}|\bm{\theta})$ is the likelihood of observing the trajectory $\bm{q}_{[1:N]}$ given $\bm{\theta}$, and $p(\bm{q}_{[1:N]})$ is a normalization factor.
@@
Typical inference algorithms maximize the likelihood with respect to the parameters $\bm{\theta}$. However, evaluating the likelihood for partially observable dynamical models is very challenging, because it is a marginalized likelihood over all possible “latent” trajectories of the actual molecular extension $\bm{x}_{[1:N]}$ [4, 11]:
{align}
p(q_[1:N]|θ) = ∫Dx_[1:N] p(q_[1:N], x_[1:N]|θ)
@@
An emerging approach to tackle this problem is simulation-based inference (SBI) [5, 10, 11], which integrates physics-based models with probabilistic deep learning to learn a surrogate model of the posterior [39], likelihood [41], or likelihood ratio [15]. The core idea is to use a simulator to generate synthetic measurements for specific parameters $\bm{\theta}$. Using deep neural networks, we can learn a surrogate model of the posterior distribution from the simulated data, providing a probabilistic mapping between parameters and data. SBI has shown success across various fields, from astrophysics [7, 45] to neuroscience [18, 29], particularly for models with complex likelihoods [28].
@@
In this work, we showcase how to leverage simulation-based inference to learn quantitative folding models from single-molecule force spectroscopy experiments by learning the Bayesian posterior from synthetic data. This framework yields accurate uncertainty estimates and quantitative predictive models, providing an efficient and robust approach for analyzing SMFS data. We applied it to the well-characterized 30R50/T4 DNA hairpin as a model system [33, 35, 31, 36], and showed that it can reconstruct the hairpin’s free energy landscape even from limited experimental data. We further demonstrated the generality of the approach by applying it to the multi-state folding of a riboswitch aptamer, highlighting its applicability to more complex systems.
@@
Single nucleic acid hairpins fold into simple structures, offering a well-defined and biologically relevant experimental framework for investigating folding mechanisms [27, 50, 6]. The end-to-end distance of the DNA serves as a convenient reaction coordinate for the folding reaction, along which the dynamics are well characterized by one-dimensional Brownian diffusion on an effective free energy landscape [33, 34].
@@
We applied our inference framework [10] to learn a quantitative folding model from constant-force measurements of a 30R50/T4 DNA hairpin. This particular hairpin has been extensively studied, with measurements of the folding dynamics spanning from microseconds to minutes and folding energy landscape reconstructions using various methods and experimental setups [53, 16, 33, 22, 30].
@@
The physical model $\mathcal{M}(\bm{\theta})$ describes both the biomolecular folding process and the measurement of the observed data. Given a measured trajectory $\bm{q}_{[1:N]}$, we aim to determine the optimal parameter set $\bm{\theta}$ that best fits the measurement by computing the posterior $p(\bm{\theta}|\bm{q}_{[1:N]}$).
@@
The harmonic-spring model $\mathcal{M}(\bm{\theta})$ provides a well-established framework for describing single-molecule force spectroscopy experiments [4, 3, 26]. It describes the coupled dynamics of the biomolecule and the measuring apparatus as a diffusive process on a two-dimensional free energy surface $G(q,x)$. Here, $q$ is the measured extension, including both the linker and the molecule, while $x$ is the (hidden) molecular extension. The free energy surface decomposes as $G(q,x)=G_{0}(x)+1/2\cdot k_{l}(q-x)^{2}$ where $G_{0}(x)$ is the intrinsic folding free energy landscape of interest. The second term describes the linker coupling to the apparatus. We assume anisotropic Brownian dynamics on this energy landscape, with diffusion coefficient $D_{x}$ for the molecule and $D_{q}$ for the apparatus. The molecular free energy profile $G_{0}(x)$ is modeled using cubic spline interpolation. The model parameters $\bm{\theta}=\{D_{q}/D_{x},k_{l},G_{0}(x_{0}),G_{0}(x_{1}),...,G_{0}(x_{N_{\textrm{node}}})\}$ are the ratio of diffusion coefficients $D_{q}/D_{x}$, the linker stiffness $k_{l}$, and $N_{\textrm{node}}$ positions of the spline nodes $G_{0}(x_{i})$ to characterize the free energy profile $G_{0}(x)$. Inferring $\bm{\theta}$ from a measured trajectory $\bm{q}_{[1:N]}$ fully characterizes biomolecular folding as diffusion on the one-dimensional landscape $G_{0}(x)$.
@@
Direct computation of the likelihood (Eq. Quantitative and Predictive Folding Models from Limited Single-Molecule Data Using Simulation-Based Inference) is extremely challenging because it requires integrating over all hidden molecular trajectories $\bm{x}_{[1:N]}$. In other words, computing the likelihood involves searching over all possible paths $\bm{x}_{[1:N]}$ consistent with the observed trajectory $\bm{q}_{[1:N]}$. In contrast, simulating a measured trajectory $\bm{q}_{[1:N]}\sim p(\bm{q}_{[1:N]}|\bm{\theta})$ is straightforward. SBI offers a general solution to perform Bayesian inference in settings where the likelihood is either intractable or prohibitively expensive, but forward simulations are feasible. SBI starts by generating prior samples $\bm{\theta}_{i}\sim p(\bm{\theta})$, which are used to simulate synthetic trajectories $\bm{q}_{[1:N],i}\sim\mathcal{M}(\bm{\theta}_{i})$ (Fig. 1 upper panel). The simulated data set $\mathcal{D}=\{(\bm{q}_{[1:N],i},\bm{\theta}_{i})\}_{i=1}^{M}$ is then used to train a neural network-based density estimator $f_{\psi}(\bm{\theta}|\bm{q}_{[1:N]})$ of parameters $\psi$ to approximate the Bayesian posterior (Fig. 1 center panel). The neural network parameters $\psi$ are optimized by minimizing the loss function $\mathcal{L}(\psi)=1/M\sum_{i}^{M}f_{\psi}(\bm{\theta}_{i}|\bm{q}_{[1:N],i})$ on $\mathcal{D}$. After training, the posterior for an experimental trajectory is obtained as $p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})\approx f_{\psi}(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$, enabling inference of the folding parameters (Fig. 1 right panel).
@@
We analyzed data from the 30R50/T4 DNA hairpin (Fig. 2A Cartoon). Our analysis used a short 2-second long constant force measurement $\bm{q}_{[1:N]}^{\textrm{exp}}$ (Fig. 2 A) that captured approximately seven transitions between the folded and unfolded states.
Figure 2: Free energy profile reconstruction. (A) Experimental time series used for inference. (B) Reconstructed free energy profile. Best estimate $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}$ (Maximum a posteriori, MAP) in red, and posterior samples covering a 68 % confidence interval as blue thin lines. The black line indicates the estimate using deconvolution. (C) Best free energy profile (MAP) estimate for 20 independent experimental time series.
@@
We used Sequential Neural Posterior Estimation (SNPE) to approximate the posterior distribution $p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$ for the experimental trajectory $\bm{q}_{[1:N]}^{\textrm{exp}}$. The trained surrogate posterior can be evaluated on the experimental data, allowing parameters to be sampled from the posterior, $\bm{\theta}^{\textrm{exp}}\sim p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$, to explore many parameter combinations consistent with the experiment. Alternatively, identifying the maximum a posteriori estimate, $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}=\textrm{argmax}_{\bm{\theta}}\ p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$, yields the best-fitting parameter set. SNPE approximates the posterior through iterative rounds of simulation and posterior density estimation. In each round, we carried out simulations with parameters drawn from the current posterior approximation. After 14 such rounds, each consisting of 2000 simulations, we obtained a surrogate posterior $p(\bm{\theta}|\bm{q}_{[1:N]}^{\textrm{exp}})$ (Fig. S1) in approximately 11 hours on a single Xeon Skylake Gold 6148 CPU.
@@
Figure 3: Diffusion coefficients and linker stiffness estimates. Posteriors obtained using 20 independent experimental time series, quantifying the inference on (A) the ratio of diffusion coefficients $D_{q}/D_{x}$, and (B) the linker stiffness $k_{l}$. Posterior marginals were obtained by sampling and histogramming. All marginals are normalized to unit area. A constant vertical offset is applied between marginals to aid visualization
@@
We validated our results by comparing our inferred free energy landscape with one obtained via deconvolution (Fig. 2B black dotted line). Deconvolution is an established method that reconstructs the free energy landscape by iteratively removing measurement noise from the experimental data (see SI) [53]. Using measurements of the biomolecule-linker system $G(q)$ and the linker only V(qx)V(q-x), deconvolution aims to find $G_{0}(x)$, that, when convolved with V(qx)V(q-x), reproduces $G(q)$. While both approaches yield closely matching profiles, confirming the reliability of our inference, deconvolution requires substantially more experimental data — approximately 20 - 100 times as many folding/unfolding transitions. This makes deconvolution more labor-intensive [53] and prone to errors arising from instrumental drift. Furthermore, deconvolution requires a separate linker characterization, where inaccuracies can introduce errors in the resulting landscape.
@@
Figure 4: Predictive checks using simulations with best-fitting parameters $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}$. (A) Trajectory $\bm{q}_{[1:N],1}$ simulated with $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},1}$ (blue) compared to experimental trajectory $\bm{q}_{[1:N],1}^{\mathrm{exp}}$ used to make the inference (red). (B) The potential of mean force estimated from 20 synthetic trajectories $\bm{q}_{[1:N],i}$ (blue) simulated with $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},1}$ compared to the potential of mean force obtained from the experimental trajectories (red). (C) Autocorrelation functions for segments of the trajectories in the folded and unfolded states, respectively, comparing experimental $\bm{q}_{[1:N],i}^{\mathrm{exp}}$ (red) and simulated trajectories $\bm{q}_{[1:N],i}$ (blue).
Additionally, we demonstrated that our estimated uncertainties are accurate by applying our inference framework to 20 independent 2-second measurements $\{\bm{q}_{[1:N],i}^{\mathrm{exp}}\}_{i=1}^{20}$ of the same DNA hairpin molecule. The results, shown in Figure 2C, yield statistically consistent free energy landscapes across all measurements. Deviations among these independently inferred profiles (red lines in Figure 2C) fall within the uncertainty bounds (blue lines in Figure 2B) determined from analyzing a single trajectory, indicating that the error model of the surrogate posterior is well-calibrated.
The model we fitted to the experimental data $\bm{q}_{[1:N]}^{\textrm{exp}}$ also obtained estimates for the ratio of diffusion coefficients $D_{q}/D_{x}$ and linker stiffness $k_{l}$. Figure 3 reports the marginal posterior distribution for $D_{q}/D_{x}$ and $k_{l}$ for 20 independent experimental trajectories. We also observed a satisfactory overlap of the marginal posterior distributions for these parameters, although larger than expected, pointing to the existence of model misspecification, in particular for the diffusion coefficients (see SI).
To further assess our model’s accuracy and predictive power, we simulated trajectories using the best-fitting parameter, i.e., $\bm{q}_{[1:N],i}\sim p(\bm{q}_{[1:N]}\mid\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},1})$. We then compared these simulated trajectories to the experimental trajectory used for parameter inference (Fig. 4A). If the model is adequate and the fit accurate, we expect that the experimental $\bm{q}_{[1:N]}^{\mathrm{exp}}$ and simulated trajectory $\bm{q}_{[1:N]}$ look very similar.
@@
Visual comparison shows excellent agreement between the simulated and experimental trajectories. To quantify this, we performed a total of 20 simulations $\{\bm{q}_{[1:N],i}\}_{i=1}^{20}$ using the best fitting parameter $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP},1}$. First, we compared the thermodynamics of the observed trajectories by inspecting the potential of mean force (PMF) $G(q)$. We computed the average PMF $G(q)$ by combining all trajectories and binning them along $q$. The simulated and experimental PMFs align well in the folded and unfolded states, though the simulated trajectories exhibit a slightly higher barrier in the transition region (Fig. 4B). The variability of the individual PMFs is shown in Fig. S4.
@@
We also compared the observed kinetics by analyzing folding and unfolding transitions (see SI). The experimental transition rate of $2.8\pm 0.3\penalty 10000\ \mathrm{s^{-1}}$ matches the simulated rate of $2.2\pm 0.2\penalty 10000\ \mathrm{s^{-1}}$, demonstrating that our model captures the underlying kinetics. This agreement between the rates demonstrates that our model successfully captures the underlying kinetics.